[1](10) The table shows the rise in the percentage of births to U.S. unmarried women from 1960 through 2009 Year | x = years Does anybody think they can solve a-d? (Descriptive Statistics)
Answer #1

Sum of X = 149
Sum of Y = 136
Mean X = 24.8333
Mean Y = 22.6667
Sum of squares (SSX) = 1700.8333
Sum of products (SP) = 1261.6667

Regression Equation = y = bX + a

b = SP/SSX = 1261.67/1700.83 = 0.74179

a = MY - bMX = 22.67 - (0.74*24.83) = 4.24547

y = 0.74179X + 4.24547

B) The slope is 0.74179 and it interprets that if you increase time by 1 year the percentage of births in US by unmarried women increase by 0.74179.

c) In year 2016 so value of x = 2016 - 1960 = 56

so y = 0.74179*56+ 4.24547 = 45.7859 = 46%

d) In year 2020 so value of x = 2020 - 1960 = 60

so y = 0.74179*60+ 4.24547 = 48.7531 = 49%

If any doubt please feel free to ask through the comments section and if no doubt please give a like. Thank You